An object of mass m0 = 4 kg travels with velocity υ0, = 10 m/s and collides with a block of mass
m2 = 1 kg, initially at rest. After the collision, the object has velocity υ03 and the block velocity υ23. Both velocities have direction to the right. If the kinetic energy loss percentage of the system object – block is 20%, calculate the velocities of the block and object after the collision.
1
Expert's answer
2020-03-20T09:36:01-0400

From the conservation of momentum:


"m_0u_0=m_0u_{03}+m_2u_{23}\to (4)(10)=4u_{03}+u_{23}"

"(1-0.2)0.5m_0u_0^2=0.5m_0u_{03}^2+0.5m_2u_{23}^2"

"2(0.8)(10)^2=2u_{03}^2+0.5((4)(10-u_{03}))^2"

"u_{03}=8\frac{m}{s}"


"u_{23}=10-8=2\frac{m}{s}"